Short Answer Question 5:

Estimate the market value per share of a firm with the following characteristics:

• Annual revenues are $775,000, cash expenses as a % of revenues of 20%, no depreciation expense, tax rate of 20%. These variables will not change over the projection period.

• The firm will earn after tax cash flows from operations for three more years.

• At the end of three years, the firm will be sold at a price estimated using a P/E ratio multiple of 35. Assume that there will be no tax effects associated with the sale of the firm.

• The firm has shares outstanding of 10.

• The firm’s WACC is 10%.
Answer #1

Free cash flow of firm for 3 years

= Revenue-expense-tax =775000- (20% of 775000)- 20% tax

=775000-155000-(20% of (775000-155000)) =620000-124000 = $496,000

Price of firm after 3 years = PE ration * earnings for year 3

=35*496000 = 17360000

Market value per share of the firm now= present value of all future Cash flows / no of shares

= (PV of 3 year free CF of $496000 + PV of terminal Value $17360000)/10 (discount rate:10% given)

=(450909.1+409917.35+372652.14+13042824.94)/10

=14276303.53/10 = $1427630.3 per share is the share value